Q)Calculate the ionic strength value for half-normal saline, containing sodium chloride

(c(NaCl) = 7.70×10-3 mol/L) and dextrose (c(dextrose) = 0.28 mol/L).

(a) 7.70×10^-3 mol/L

(b) 0.0154 mol/L

(c) 0.1440 mol/L

(d) 0.2877 mol/L


Q)Calculate the boiling point of hypertonic saline with mass fraction of sodium chloride

7.0%. (Kb (H2O) = 0.52 K × kg × mol-1).

(a) 1.35 K

(b) 101.35 C

(c) 0.67 K

(d) 100.67 C


Q)Sodium lauryl sulfate (SLS) is used as a surfactant in many hygiene products, in

particular in toothpastes. Calculate the surface activity of SLS at 20 C if known that surface

tension of its solution with 6.6 mmol×dm-3 concentration value is equal to 47 mN×m-1,

and with 8.3 mmol×dm-3 concentration value is equal to 43 mN×m-1.

(a) 0.43

(b) -0.43

(c) - 2.35

(d) 2.35


Calculate the pH value of acetate buffer (pKa = 4.76) prepared by mixing of 9.0 mL of

sodium acetate solution with 1.0 mL of acetic acid solution (c(CH3COOH) = c(CH3COONa)).

Determine the range of buffer action for the acetate buffer.

(a) 3.81; 3.81(+/-)1

(b) 3.81; 4.76(+/-)1

(c) 5.71; 5.71(+/-)1

(d) 5.71; 4.76(+/-)1


1
Expert's answer
2021-01-18T01:54:48-0500

Question#1 A

Question#2


"\Delta T = iEC_m"

"i = 2"

"C_m = {m \over Mm(H_2O)} = {7 \over 58.5*0.093} = 1.29"

"\Delta T = 2*0.52*1.29 = 1.3416 K"

Answer: B 101.35 oC